OpenStudy (anonymous):

multiply and simplify type coefficient as integer or decimals (0.1x^4 * 0.4) = .04 is this correct

OpenStudy (anonymous):

no, not if it is \[0.1x^4\times 0.4\]

OpenStudy (anonymous):

oh sorry. you are just asked for the coefficient. then you are correct. \[0.1\times 0.4=.04\]

OpenStudy (anonymous):

satelliet I posted wrong it is (0.1^4 * 0.4x^6) = 0.1 * 0.4 = .04 ans is this correct

OpenStudy (anonymous):

0.1x^4 sorry

OpenStudy (anonymous):

ok was it \[0.1x^4 \times 0.4x^6\]?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok if you are just asked for the coefficient you are right: .04 is the answer

OpenStudy (anonymous):

satellite is this correct simplify assume z = with line thru equal sign 0 (4z^-3)^3 = 64/z^9 ans is this correct